# Generative Unfolding of Jets and Their Substructure

### v1

@@ 1 Introduction
Experiments at the LHC observe particle interactions at high energy to study the nature of matter and forces. However, detectors have finite resolution and acceptance. Measuring physical parameters requires including these detector effects in predictions, which can be compared to data. Because such effects are expensive to model and require access to proprietary software, an alternative is the reverse, to remove the impact of detector, unfolding the data, which allows direct comparison with theoretical predictions and across experiments. The measured spectra are defined at particle level, the phase space of the stable particles before detector effects are modeled. Traditional unfolding methods require binning of the data and are limited to a small number of dimensions [1]. Machine-learning (ML)-based [2] unfolding techniques have overcome these limitations, offering un-binned unfolding in many simultaneous dimensions. Two classes of ML unfolding have emerged, discriminative and generative.
@@
Discriminative unfolding algorithms train classifiers to re-weight simulated events at particle level to produce un-binned unfolded spectra [3, 4, 5, 6, 7, 8] and have recently been used in several measurements [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. They assume that the sample of simulated events at particle level has sufficient support to construct the particle level data distribution. In this case, they are known to be accurate even in high- and variable-dimensional unfolding tasks, such as unfolding the entire substructure of high-multiplicity jets [20].
@@
Generative algorithms learn a mapping between detector-level events and particle-level events, formalized as a posterior density conditional on observed detector-level events to allow sampling from the truth-level distributions [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Generative algorithms produce unweighted events with few assumptions about the support of the simulated events and in the presence of narrow phase-space structures. However, they require solving the challenging generative ML task of modeling the full particle-level distribution. They have been shown to produce excellent accuracy for fixed-dimensional phase spaces of several dimensions [35, 33], and have also shown promise for variable-dimensional spaces [30]. However, no generative unfolding algorithm has been shown to scale to high-dimensional unfolding.
@@
Although many measurements only require a few dimensions for the final comparison, the inclusion of additional dimensions on which the detector response depends gives higher accuracy [36]. Many measurements, such as those using energy-energy correlators [37, 38, 39] are best approached by unfolding every final state particle. Unfolding without dimensionality reduction also preserves the maximum amount of information for future analysis. If the unfolding is differential in the kinematics of every final-state particle, any observable can be easily measured post-hoc, allowing reinterpretation for unforeseen use cases.
@@
This paper describes a new generative method capable of unfolding several hundred dimensions, allowing for unfolding of high-multiplicity jets. It makes use of conditional flow matching (CFM) generative networks [40, 41, 42, 43] with and without a Lorentz-equivariant architecture [44, 45] to solve three generative tasks. First, a multiplicity prediction network provides an estimate of the number of particles contained in the particle-level jet. Second, the jet kinematics are predicted and used to provide a physically motivated rescaling of the kinematics of the jet constituents. Finally, the full-dimensional phase space of the jet constituents conditioned on the predicted multiplicity and the jet kinematics is generated. The method accurately unfolds the full substructure of jets in two benchmark datasets, one containing light quark and gluon initiated jets produced in association with a $Z$-boson, and one containing jets produced by the hadronic decay of boosted top quarks.
@@ 2 Generative Unfolding
Note that the generative model always depends on a simulation prior $p(x_{\text{part}})$. In practical applications, iterative methods [26, 46] or specific setups [33, 31] can be used to remove a potential bias arising from the simulation prior.
Current state-of-the-art generative unfolding uses conditional flow matching (CFM) [40, 41, 42, 43]. It describes the transformation from a latent space $p_{\text{latent}}(z)$ to phase space $p(x_{\text{part}})$ as a time-dependent diffusion process between $t\in[0,1]$
@@
Conditional flow matching introduces conditional target trajectories [41, 42] that interpolate linearly between the latent space and phase space events
@@
An important strength of CFMs is that the velocity can be learned by any network architecture. In particle physics, encoding the symmetry structure of the phase space greatly improves the network accuracy [47, 48, 44, 45, 49, 50, 51, 52]. For this reason, the velocity is best learned by permutation equivariant transformer architectures [53], or transformers with an additional Lorentz-equivariant constraint [44, 45] like L-GATr, as illustrated in Figure 1.
@@ 3.2 Jet unfolding
with each component standardized. The network receives the multiplicity $N_{\text{part}}$ as well as the detector-level jet kinematics $J_{\text{reco}}$ as conditions. Though the detector level jet kinematics can be sensitive to “hidden-variables” which produce a non-universal forward simulation [34], in this study it is sufficient to use only these inputs instead of the full detector level constituent kinematics information $x_{\text{reco}}$. Additionally, alternative parametrizations of the jet kinematics $J_{\text{part}}$ can help the network to learn task-specific features. See Section 4.
@@
For the L-GATr, the condition is embedded using an encoder similar to the multiplicity unfolding, but now including only the four-momentum of the detector-level jet $J_{\text{reco}}$ and constituent multiplicity $N_{\text{reco}}$ instead of the full constituent information $x_{\text{reco}}$. The beam and time directions are also included to allow for symmetry breaking. The decoder for the velocity prediction takes the noised particle-level jet four-momentum, parametrized in rectilinear coordinates as $(E_{J},p_{x,J},p_{y,J},p_{z,J})$ as input, together with the time embedded in a scalar channel, and returns a four-velocity $(v_{E_{J}},v_{p_{x,J}},v_{p_{y,J}},v_{p_{z,J}})$. The linear CFM trajectories are defined in the $(\log p_{T,J},{}_{J},{}_{J},\log m_{J}^{2})$ parametrization, so the particle-level jet $J_{\text{part}}$ is transformed to rectilinear coordinates before input to the flow matching network. After extracting the four-velocity and multiplying by the inverse Jacobian of the forward transformation, the velocity field $(v_{\log p_{T,J}},v_{{}_{J}}v_{{}_{J}},v_{\log m_{J}^{2}})$ is obtained. To avoid numerical instabilities from large Jacobians, after the transformation, the $v_{\log p_{T,J}}$ and $v_{\log m_{J}^{2}}$ components are over-written with additional Lorentz scalar outputs of the L-GATr network, while the $v_{{}_{J}}$ and $v_{{}_{J}}$ components are kept [44].
@@ 5 Unfolding boosted top jets
The precision-generative unfolding method is additionally applied to the task of unfolding the substructure of jets originating from the all-hadronic decay of boosted top quarks. Boosted hadronically decaying top jets have become a standard analysis object at the LHC [60, 61] and can even be used to extract the top mass [62, 39, 33].
@@
is generated with Madgraph 5 [63] assuming a top mass of $m_{t}=173$ GeV. Parton showering and hadronization are simulated with Pythia 8.313 [54] and the detector response is simulated with Delphes 3.5.0 [55] using the default CMS card. After reconstruction, detector-level and particle-level jets are defined via the anti-$k_{\text{T}}$ algorithm [56] with $R=1.2$ using FastJet [57]. Each event is required to have exactly two jets with $p_{\text{T}}^{J}>10$ GeV and exactly one lepton satisfying $p_{\text{T}}^{\ell}>60$ GeV and $|{}^{\ell}|<2.4$. In addition, each event must have exactly two $b$-tagged sub-jets, clustered with anti-$k_{\text{T}}$ and $R=0.4$. The jet with the largest angular distance to the lepton is tagged as the top jet. The top jet is additionally required to satisfy
@@ 6 Conclusion
The performance shown in this study marks the state of the art for generative unfolding. The main limitation of this study has been the size of the transformer networks, their training time, and the size of the training dataset. For classifier-based unfolding network ensembles have been shown to lead to significant performance increase. As shown for L-GATr jet tagging [44], larger pre-trained transformers should lead to further improvements of the accuracy.

### v2

@@ 1 Introduction
Experiments at the LHC observe particle interactions at high energy to study the nature of matter and forces. However, detectors have finite resolution and acceptance. Measuring physical parameters requires including these detector effects in predictions, which can be compared to data. Because such effects are expensive to model and require access to proprietary software, an alternative is the reverse, to remove the impact of detector, unfolding the data, which allows direct comparison with theoretical predictions and across experiments. The measured spectra are defined at particle level, the phase space of the stable particles before detector effects are modeled. Traditional unfolding methods require binning of the data and are limited to a small number of dimensions [Cowan:2002in]. Machine-learning (ML)-based [Plehn:2022ftl] unfolding techniques have overcome these limitations, offering un-binned unfolding in many simultaneous dimensions. Two classes of ML unfolding have emerged, discriminative and generative.
@@
Discriminative unfolding algorithms train classifiers to re-weight simulated events at particle level to produce un-binned unfolded spectra [Andreassen:2019cjw, Chan:2023tbf, Desai:2024yft, Zhu:2024drd, Falcao:2025jom, Canelli:2025ybb] and have recently been used in several measurements [H1:2021wkz, H1:2023fzk, H1:2024mox, LHCb:2022rky, Huang:2025ziq, Pani:2024mgy, Song:2023sxb, ATLAS:2024rpl, ATLAS:2025qtv, CMS:2025sws, Badea:2025wzd]. They assume that the sample of simulated events at particle level has sufficient support to construct the particle level data distribution. In this case, they are known to be accurate even in high- and variable-dimensional unfolding tasks, such as unfolding the entire substructure of high-multiplicity jets [Komiske:2022vxg].
@@
Generative algorithms learn a mapping between detector-level events and particle-level events, formalized as a posterior density conditional on observed detector-level events to allow sampling from the truth-level distributions [Datta:2018mwd, Bellagente:2019uyp, Bellagente:2020piv, Vandegar:2020yvw, Howard:2021pos, Backes:2022sph, Diefenbacher:2023wec, Ackerschott:2023nax, Shmakov:2023kjj, Shmakov:2024gkd, Pazos:2024nfe, Butter:2024vbx, Favaro:2025psi, Butter:2025via]. Generative algorithms produce unweighted events and could have complementary strengths and weaknesses compared to discriminative methods. However, they require solving the challenging generative ML task of modeling the full particle-level distribution. They have been shown to produce excellent accuracy for fixed-dimensional phase spaces of several dimensions [Huetsch:2024quz, Favaro:2025psi], and have also shown promise for variable-dimensional spaces [Shmakov:2024gkd]. However, no generative unfolding algorithm has been shown to scale to high-dimensional unfolding.
@@
Although many measurements only require a few dimensions for the final comparison, the inclusion of additional dimensions on which the detector response depends gives higher accuracy [Desai:2025mpy]. Many measurements, such as those using energy-energy correlators [Lee:2022uwt, Moult:2025nhu, Holguin:2024tkz] are best approached by unfolding every final state particle. Unfolding without dimensionality reduction also preserves the maximum amount of information for future analysis. If the unfolding is differential in the kinematics of every final-state particle, any observable can be easily measured post-hoc, allowing reinterpretation for unforeseen use cases.
@@
This paper describes a new generative method capable of unfolding several hundred dimensions, allowing for unfolding of high-multiplicity jets. It makes use of conditional flow matching (CFM) generative networks [chen2018neural, lipman2023flowmatchinggenerativemodeling, albergo2023stochastic, Butter:2023fov] with and without a Lorentz-equivariant architecture [Brehmer:2024yqw, Favaro:2025pgz] to solve three generative tasks. First, a multiplicity prediction network provides an estimate of the number of particles contained in the particle-level jet. Second, the jet kinematics are predicted and used to provide a physically motivated rescaling of the kinematics of the jet constituents. Finally, the full-dimensional phase space of the jet constituents conditioned on the predicted multiplicity and the jet kinematics is generated. The method accurately unfolds the full substructure of jets in two benchmark datasets, one containing light quark and gluon initiated jets produced in association with a $Z$-boson, and one containing jets produced by the hadronic decay of boosted top quarks.
@@ 2 Generative Unfolding
Note that the generative model always depends on a simulation prior $p(x_{\text{part}})$. In practical applications, iterative methods [Backes:2022sph, Butter:2025mek] or specific setups [Favaro:2025psi, Pazos:2024nfe] can be used to remove a potential bias arising from the simulation prior.
Current state-of-the-art generative unfolding uses conditional flow matching (CFM) [chen2018neural, lipman2023flowmatchinggenerativemodeling, albergo2023stochastic, Butter:2023fov]. It describes the transformation from a latent space $p_{\text{latent}}(z)$ to phase space $p(x_{\text{part}})$ as a time-dependent diffusion process between $t\in[0,1]$
@@
Conditional flow matching introduces conditional target trajectories [lipman2023flowmatchinggenerativemodeling, albergo2023stochastic] that interpolate linearly between the latent space and phase space events
@@
An important strength of CFMs is that the velocity can be learned by any network architecture. In particle physics, encoding the symmetry structure of the phase space greatly improves the network accuracy [Gong:2022lye, Bogatskiy:2022czk, Brehmer:2024yqw, Favaro:2025pgz, Witkowski:2023xmx, Nabat:2024nce, Breso:2024jlt, Bahl:2024gyt]. For this reason, the velocity is best learned by permutation equivariant transformer architectures [Heimel:2023mvw], or transformers with an additional Lorentz-equivariant constraint [Brehmer:2024yqw, Favaro:2025pgz] like L-GATr, as illustrated in Figure 1.
@@ 3.2 Jet unfolding
with each component standardized. The network receives the multiplicity $N_{\text{part}}$ as well as the detector-level jet kinematics $J_{\text{reco}}$ as conditions. Though the detector level jet kinematics can be sensitive to “hidden-variables” which produce a non-universal forward simulation [Butter:2025via], in this study it is sufficient to use only these inputs instead of the full detector level constituent kinematics information $x_{\text{reco}}$. Additionally, alternative parametrizations of the jet kinematics $J_{\text{part}}$ can help the network to learn task-specific features. See Section 4.
@@
For the L-GATr, the condition is embedded using an encoder similar to the multiplicity unfolding, but now including only the four-momentum of the detector-level jet $J_{\text{reco}}$ and constituent multiplicity $N_{\text{reco}}$ instead of the full constituent information $x_{\text{reco}}$. The beam and time directions are also included to allow for symmetry breaking. The decoder for the velocity prediction takes the noised particle-level jet four-momentum, parametrized in rectilinear coordinates as $(E_{J},p_{x,J},p_{y,J},p_{z,J})$ as input, together with the time embedded in a scalar channel, and returns a four-velocity $(v_{E_{J}},v_{p_{x,J}},v_{p_{y,J}},v_{p_{z,J}})$. The linear CFM trajectories are defined in the $(\log p_{T,J},{}_{J},{}_{J},\log m_{J}^{2})$ parametrization, so the particle-level jet $J_{\text{part}}$ is transformed to rectilinear coordinates before input to the flow matching network. After extracting the four-velocity and multiplying by the inverse Jacobian of the forward transformation, the velocity field $(v_{\log p_{T,J}},v_{{}_{J}}v_{{}_{J}},v_{\log m_{J}^{2}})$ is obtained. To avoid numerical instabilities from large Jacobians, after the transformation, the $v_{\log p_{T,J}}$ and $v_{\log m_{J}^{2}}$ components are over-written with additional Lorentz scalar outputs of the L-GATr network, while the $v_{{}_{J}}$ and $v_{{}_{J}}$ components are kept [Brehmer:2024yqw].
@@ 5 Unfolding boosted top jets
The precision-generative unfolding method is additionally applied to the task of unfolding the substructure of jets originating from the all-hadronic decay of boosted top quarks. Boosted hadronically decaying top jets have become a standard analysis object at the LHC [Plehn:2009rk, Plehn:2011tg] and can even be used to extract the top mass [CMS:2022kqg, Holguin:2024tkz, Favaro:2025psi].
@@
is generated with Madgraph 5 [Alwall:2011uj] assuming a top mass of $m_{t}=173$ GeV. Parton showering and hadronization are simulated with Pythia 8.313 [Sjostrand:2014zea] and the detector response is simulated with Delphes 3.5.0 [deFavereau:2013fsa] using the default CMS card. After reconstruction, detector-level and particle-level jets are defined via the anti-$k_{\text{T}}$ algorithm [Cacciari:2008gp] with $R=1.2$ using FastJet [Cacciari:2011ma]. Each event is required to have exactly two jets with $p_{\text{T}}^{J}>10$ GeV and exactly one lepton satisfying $p_{\text{T}}^{\ell}>60$ GeV and $|{}^{\ell}|<2.4$. In addition, each event must have exactly two $b$-tagged sub-jets, clustered with anti-$k_{\text{T}}$ and $R=0.4$. The jet with the largest angular distance to the lepton is tagged as the top jet. The top jet is additionally required to satisfy
@@ 6 Conclusion
The performance shown in this study marks the state of the art for generative unfolding. The main limitation of this study has been the size of the transformer networks, their training time, and the size of the training dataset. For classifier-based unfolding network ensembles have been shown to lead to significant performance increase. As shown for L-GATr jet tagging [Brehmer:2024yqw], larger pre-trained transformers should lead to further improvements of the accuracy.